The specific heat of a certain type of cooking oil is 0.418 cal/(g·°C). How much heat energy is needed to raise the temperature of 1.069 kg of this oil from 23.0 °C to 60.0 °C?
Answer #1

Given:
m = 1069 g
C = 0.418 cal/g.oC
Ti = 23 oC
Tf = 60 oC
use:
Q = m*C*(Tf-Ti)
Q = 1069.0*0.418*(60.0-23.0)
Q = 1.653*10^4 cal
Answer: 1.65*10^4 cal